Given: d = rt, where "d" is distance, "r" is rate, and "t" is time.
A plane flies a distance of 2,320 miles in 5.8 hours. What is its
average speed?

Answer by jim_thompson5910(35256)

d = r*t


2320 = r*5.8


2320/5.8 = r


400 = r


r = 400


The average speed is 400 mph